OpenStudy (anonymous):

WILL MEDAL!!!!!!!!!!! Laurie is trying to stay within 10 feet of her current diving depth of –30 feet (with regard to sea level) so that the light is still good but she can be close to the sea life during her scuba dive. Which two equations can be used to find the minimum and maximum depths Laurie wants to stay between? x + 30 = 10 and x + 30 = –10 x – 30 = 10 and x – 30 = –10 x + 10 = 30 and x + 10 = –30 x – 10 = 30 and x – 10 = –30

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

the second one

OpenStudy (aum):

Let x be the depth. The difference between x and the depth of -30 must be less than or equal to 10. That is, |x - (-30)| <= 10 or |x + 30| <= 10 The minimum and maximum depth is when |x + 30| = 10 which implies x + 30 = -10 and x + 30 = 10